### v2

# See, Imagine, Plan: Discovering and Hallucinating Tasks from a Single Image

Abstract
Humans can not only recognize and understand the world in its current state but also envision future scenarios that extend beyond immediate perception. To resemble this profound human capacity, we introduce zero-shot task hallucination—given a single RGB image of any scene comprising unknown environments and objects, our model can identify potential tasks and imagine their execution in a vivid narrative, realized as a video. We develop a modular pipeline that progressively enhances scene decomposition, comprehension, and reconstruction, incorporating VLM for dynamic interaction and 3D motion planning for object trajectories. Our model can discover diverse tasks, with the generated task videos demonstrating realistic and compelling visual outcomes that are understandable by both machines and humans. All code, data, and additional videos are available at our project page: https://dannymcy.github.io/zeroshot_task_hallucination/
“A rock pile ceases to be a rock pile the moment a single man contemplates it, bearing within him the image of a cathedral.” —Antoine de Saint-Exupéry
Figure 1: We present a model for zero-shot task hallucination. With a single RGB image of any scene comprising unknown environments and objects, we identify potential tasks (task discovery) and imagine their execution (manipulation) in a vivid narrative, realized as a video. We also provide the trajectories of the executed tasks.
1 Introduction
The theory of mental time travel [59] highlights the human ability to recall past knowledge and anticipate future scenarios through visual perceptions. For example, the sight of a fallen chair immediately prompts in our mind an imagined rehearsal to straighten it back up. Equipping intelligent agents with this level of imaginative foresight and the ability to visualize such imagination could open up exciting possibilities, such as serving interactive tutors, offering step-by-step visual guidance on tasks, or teaching robots to autonomously discover and undertake tasks in new environments.
In this work, we aim to resemble this profound human capability to see, imagine, and plan, by utilizing the recent advancements of large foundation models [7, 47]. We introduce the problem of zero-shot task hallucination. Given a single RGB image of any scene comprising unknown environments and objects, our objective is to identify potential tasks (task discovery) and imagine their execution (manipulation) in a vivid narrative, realized as a video, as shown in Fig 1. The proposed tasks within the 3D scenes should be diverse yet reasonable. Moreover, the generated task videos should be geometric-aware and photo-realistic. The object motion should be feasible, fine-grained, and consistent to ensure the task execution is interpretable by both humans and machines.
The challenge of zero-shot task hallucination is threefold. First, the model must encompass strong capabilities of scene understanding. This includes not only the ability to recognize arbitrary objects from diverse scene settings, but also the understanding of interrelations between objects to propose reasonable tasks. Second, the model must have a 3D understanding of the scene so that the planned task execution follows the given spatial constraints. Finally, this hallucinated task execution should be presented by a human-interpretable fashion (e.g., a video). The closest existing approaches to addressing these challenges lie in video generation and 3D-aware image editing fields. Yet, no current solution seamlessly combines high-quality task execution visualization with geometric awareness. State-of-the-art (SOTA) video generation models [52, 32] are confined to 2D pixel space. They show satisfactory performance only in short sequences and struggle with the precision needed for detailed, prolonged manipulation tasks. Leading 3D-aware image editing models [18, 48] fail to maintain smooth transitions and consistency across interpolated frames. Moreover, both approaches lack the ability to understand 3D scenes and the capability to process detailed, geometry-aware prompts (e.g., covering the pot with the pot lid), which are needed for manipulation tasks.
Our solution addresses these challenges with a modular pipeline that progressively enhances scene decomposition, comprehension, and reconstruction, incorporating Vision-Language Model (VLM) for dynamic interaction and 3D motion planning for object trajectories, producing geometric-aware task videos. To understand the image scene, we use VLM to identify interactive objects and propose context-dependent tasks in a role-play manner, complemented by language-guided segmentation and repainting models to obtain occlusion-free object masks. Elevating 2D understanding to 3D, we use depth estimation and single-view 3D reconstruction models to generate a semi-reconstructed 3D scene, with the full 3D representation of foreground objects and the background as a plane. With the reconstructed 3D scene, we introduce a novel axes-constrained 3D planning approach that enables VLM to plan the motion of objects for given tasks by specifying waypoints.
Through the combination of traditional path planning algorithms, our model generates complete, feasible, and natural trajectories from merely a single image observation. With the entire framework being fully modularized, each component can be easily replaced with the latest improvements within its specific domain.
In summary, our main contributions are threefold:
  • $\bullet$
    We study zero-shot task hallucination—the capability for models to discover and propose possible tasks and plans of execution given a single image.
  • $\bullet$
    We devise a plug-and-play framework that leverages large pretrained VLM and 3D reconstruction model, combining with traditional path planning algorithms to provide geometric-aware trajectories for diverse tasks.
  • $\bullet$
    We show our model can convert these task plans into human-interpretable formats such as videos for various potential applications, supported by extensive experiments.
2 Related Work
3D-Aware Image Editing   The introduction of diffusion models [24, 56] has created a plethora of high-performing aesthetic text-to-image generative models [24, 54, 3, 53, 50]. Various editing methods soon followed to bring more controls over the images generated. [6, 23, 49] enables edits to real-world images with additional words, while [19] allows fine-grained controls over the pixel space.
Several works also explored precise edits over the 3D space (e.g., object rotation, translation, illumination, non-rigid shape changes) [44, 12, 57, 48]. Zero-1-to-3 [40], ViewNETI [8], and Continuous 3D Words [12] enable rotations of an object in the image, but are generally single-object centric. 3DIT uses large-scale 3D datasets [18] to simulate and learn scene arrangements, while Diffusion Handles [48] elevates the image editing space to 3D via depth estimation.
Inspired by these works, and in light of the complicated tasks that could be discovered in our pipeline, and the recent advancements of many generalizable single-image object reconstruction techniques [39, 38, 2], we propose to explicitly reconstruct part of the scene to perform edits and rerender back to 2D.
Task Discovery   Previous research on robotic agent targeted predefined tasks [27, 42] or environments [20, 61]. Very recently, a few works [66, 1] proposed the concept of task discovery/proposal. RoboGen [66] presented the idea of generative simulation, in which they leveraged foundation models to generate robotic learning data by configuring new scenes and proposing robotic tasks in physics engines. AutoRT [1] observed and proposed tasks with a large fleet of robots. Nevertheless, both of these works made the underlying assumption that the scenes either created or observed are suitable for operation by certain robot types. Our work diverges from these constraints and is applicable to all real scenes. Moreover, our method requires only a single RGB image as observation/input.
Spatial Reasoning and Planning with VLMs   With the internet-scale training data and the birth of LLMs [7, 47], VLMs are encompassed with reasoning and scene understanding capabilities [11, 31, 37]. Several works build from the foundation and establish datasets to help with spatial reasoning/understanding [28, 36, 44]. As a concurrent work, Spatial VLM [10] enhanced standard VLMs with the ability to reason and answer questions requiring spatial understanding. In this work, we take one step further—asking VLMs to not only reason about the spatial information, but also plan the trajectory for tasks/actions in 3D space.
3 Method
Figure 2: The proposed framework. Our plug-and-play framework is fully modularized and zero-shot. Each module can be easily replaced with the latest improvements. Exact prompts for VLM are presented in Appendix. J.
Given a single RGB image $I\in\mathbb{R}^{H\times W\times 3}$ of a scene with $K$ unknown objects, our goal is to discover $Q$ tasks suitable for execution by a robotic/human hand, and generate $Q$ task videos, each visually depicting the execution of these tasks.
Our fully modularized framework progressively enhances scene decomposition, comprehension, and reconstruction, incorporating VLM for dynamic interaction and 3D motion planning for object trajectories, producing geometric-aware task videos (Fig. 2). It begins by understanding 2D image scene (Sec. 3.1). VLM identifies interactive objects and proposes context-dependent tasks through role-play, complemented by a two-step segmentation to acquire occlusion-free object masks with language-guided segmentation and repainting models. We then reconstruct the 3D scene in a semi-reconstructed fashion leveraging single-view 3D reconstruction and depth estimation models (Sec. 3.2). Our method for camera pose and object poses and sizes estimation does not require camera intrinsic matrix and uses only normalized depth, thus can be generalized to any input image. With the reconstructed 3D scene, we conduct in-context learning to let VLM understand the 3D spatial context. Subsequently, we introduce a novel axes-constrained 3D planning approach that enables VLM to plan the object motion based on the proposed tasks by specifying waypoints (Sec. 3.3). To generate complete, feasible, and natural trajectories, we use traditional robotics path planning and optimization methods. Finally, we execute the tasks within our reconstructed 3D scene following the planned paths, and render as videos.
3.1 Understanding Objects and Context of 2D Image Scene
3.1.1 Interactive Objects Identification
We start with querying VLM for intractable objects identification given $I$. We explicitly ask VLM to identify and name useful objects in term of affordances. Note that we exclude background objects (e.g., window, ceiling, stair), as they tend to be fixed and typically offer limited interaction opportunities. We also prompt VLM to provide concise descriptions, capturing the distinct color and texture of each identified object. This step is vital for differentiating between items of similar categories and appearances.
As a concrete example, given the left image of Fig. 2, VLM outputs:
‘‘object 0: laptop of color black and gold with a smooth texture.
object 1: camera of color black with a textured grip. …’’
3.1.2 Occlusion-Free Object 2D Masks and Background Inpainting
Occlusion-Free Object 2D Masks   The concise descriptions of identified objects are used as input text prompts for a language-guided segmentation model, enabling the acquisition of $K$ segmentation masks $\{M_{k}^{occ}\}_{k=1}^{K}$, with each mask corresponding to a unique object.
However, an object $i\in[1,K]$ may be occluded by other object(s), leading to an incomplete mask $M_{i}^{occ}$. To resolve this, we create an inpainting mask, $M_{i}^{inp}$, for each object, in which all objects except the one itself are removed and replaced with white pixels. The inpainted masks are again fed to the language-guided segmentation model along with input text prompts such that occlusion-free object masks, $\{M_{k}^{of}\}_{k=1}^{K}$, are obtained. This two-step segmentation process for object $i$ can be formulated as:
$M_{i}^{occ}=\text{seg}(I,\tau_{i}),\quad M_{i}^{of}=\text{seg}(\text{inpaint}(M_{i}^{inp}),\tau_{i}),$ (1)
where seg denotes the language-guided segmentation and $\tau_{i}$ denotes the description of object $i$. In practice, to cleanly remove objects without residual fragments for inpainting, we apply dilation to and expand the white areas.
Background Inpainting   We also remove and replace all objects with dilated white pixels, and inpaint to acquire the inpainted background $I_{bg}$.
3.1.3 Task Proposal
We query VLM to propose meaningful, diverse tasks, each with a one-sentence task description. Instead of directly querying VLM for task proposal, we employ a hybrid approach that integrates role-play and object-based initialization. In the role-play scenario, we prompt VLM to envision itself as a robotic/human hand working in the scene to perform household tasks. For the object-based initialization, we guide VLM to sequentially focus on each identified object within the scene. When the scene contains more than one identified objects, VLM is instructed to suggest two tasks emphasizing interactions between the manipulating object and any of the detected objects, and an additional task focused solely on the manipulating object. If only one object is detected, VLM is directed to propose a task involving just that object. This strategy guarantees a broad spectrum of task suggestions, ensuring comprehensive object engagement.
To further tailor the task proposals, we impose specific constraints, directing VLM to consider the practical affordances of objects while encouraging creative assumptions (e.g., a bowl’s capacity to hold water) and potential interactions (e.g., transferring water from a cup into a bowl). Additionally, we delineate clear boundaries by excluding tasks that entail the assembly or disassembly of objects, functionality tests, or the involvement of imaginary objects, thereby focusing on feasible and meaningful tasks.
As a concrete example, given the image on the left of Fig. 2, with the manipulating object to be the red can, VLM will propose the following tasks:
‘‘Task name: Can to Bowl Transfer
Description: Pick up the can and pour its contents into the bowl.
Task name: Can Relocation
Description: Pick up the can and place it inside the bowl.
Task name: Can Rotation
Description: Rotate the can 90 degrees on its vertical axis. …’’
3.2 Reconstructing and Understanding 3D Image Scene
Figure 3: Our method of partial 3D scene reconstruction. (a) full representation of foreground objects and background as a plane, (b) the reconstructed scene from the active camera view, (c) the original input image.
Our method for partial reconstruction of the 3D scene focuses on full representation of foreground objects while simplifying the inpainted background as a plane. The reconstructed scene is captured by a pinhole camera in the 3D environment, as shown in Fig. 3(a). Given that our framework is designed to accommodate input images captured by users using arbitrary cameras, we make the assumption that the camera intrinsic matrix is unknown. Our partially reconstructed scene demonstrates high visual alignment with the input image, see Fig. 3(b) and (c).
3.2.1 Single-View 3D Object Reconstruction and Depth Estimation
The first step of reconstructing the 3D image scene involves reconstructing 3D object models and estimating depth. We feed occlusion-free object masks, $\{M_{k}^{of}\}_{k=1}^{K}$, to an off-the-shelf single-view 3D reconstruction model to acquire object 3D models, $\{O_{k}\}_{k=1}^{K}$. We also feed the original map $I$ to an off-the-shelf depth estimation model to acquire depth image, denoted as $I_{dep}$.
3.2.2 Camera 6D Pose Estimation and Object Scale Initialization
Camera 6D Pose Estimation   Single-view 3D reconstruction model reconstructs mesh $O_{i}$ of object $i$ at the 3D origin, in the coordinate frame set by the input mask $M_{i}^{of}$, and captures $O_{i}$ by a pinhole camera. This camera, with 6D pose $P_{c}=[R_{c}\mid t_{c}^{origin}]$, captures $O_{i}$’s canonical pose within the image. Thus, we can restore all objects’ canonical poses across all images by identifying $P_{c}$.
We use One-2-3-45++ [38], which provides the camera pose. For models without this information, we develop an efficient method to determine the pose by comparing object masks with rendered 3D model templates, inspired from matching-based 6D pose estimation works [34, 46, 9, 41]. Details in Appendix. G.
For pinhole camera to focus on the background plane center within the reconstructed 3D scene, we backproject the 2D center of the inpainted background $I_{bg}$ using a camera intrinsic matrix $C_{\text{Mat}}$, and adjust the translation vector $t_{c}^{origin}$ in the camera pose $P_{c}$. $C_{\text{Mat}}$ is arbitrary, with the only condition of principal point coordinates, $c_{x}$ and $c_{y}$, set to image $I$ center. Focal lengths $f_{x}$ and $f_{y}$ are set to $\0\epsilon\in\mathbb{R}\backslash\{0\}$ because backprojecting the image center results in $u=c_{x}$ and $v=c_{y}$. Thus, our camera intrinsic matrix and adjusted, final camera pose are:
[…]
Setting Background Plane   The 6D pose of the background plane $O_{bg}$ (visually identical to $I_{bg}$), denoted as $P_{bg}=[R_{bg}\mid t_{bg}]$, matches the camera’s rotation $R_{c}$ as its orientation aligns with the camera. The position $t_{bg}$ is shifted along the camera’s directional vector ($v_{c}=-R_{c}\times[0,0,1]$), ensuring fitting precisely within the camera’s view based on the camera’s focal length and sensor width.
Object Scale Initialization   3D reconstruction models typically standardize objects within a normalized sphere [71, 45, 13, 39, 38]. To set the initial scale for $O_{i}$, we multiply the contour length of its occlusion-free 2D mask $M_{i}^{of}$ by a predefined scale factor $\delta_{init}$, tailored to model’s normalized space dimensions:
$S_{i}^{init}=\text{arcLength}(\text{findContour}(M_{i}^{of}))\times\delta_{init}.$ (3)
However, this method overlooks objects’ depth, inaccurately assuming uniform distance from the camera (e.g., moving an apple from close to the camera to a distant corner reduces its contour length). To correct this, we refine the object scales after determining their 6D poses, detailed in the subsequent Sec. 3.2.3.
3.2.3 Object 6D Pose Estimation and Object Scale Calibration
Object 6D Pose Estimation   Our goal is to position object 3D models into the reconstructed 3D scene without visual discrepancies and ensure accurate depth. Because a significant portion of depth estimation model [69, 51, 4, 5] is trained on depth datasets with depth data determined by sensors [21, 55] and stereo matching [70, 15, 63, 64, 67], we assume that the predicted normalized depth is the perpendicular distance to the camera plane, instead of a straight line from the object to the camera lens [30]. We start by determining the 2D centers of each occlusion-free object mask, $\{M_{k}^{of}\}_{k=1}^{K}$, using center of mass. Then, we calculate the corresponding object 3D centers on the background plane $O_{bg}$, denoted as $\{t_{k}^{bg}\}_{k=1}^{K}$. This process involves initially mapping the 2D centers onto $O_{bg}$ upon its importation into the scene and then transforming these points by multiplying the background plane’s pose $P_{bg}$. Similarly, denote the camera plane as $O_{c}$ with pose $P_{c}=[R_{c}\mid t_{c}]$ and each object 3D center on $O_{c}$ as $\{t_{k}^{c}\}_{k=1}^{K}$. For an object $i\in[1,K]$, we perform raycasting from $t_{i}^{c}$ on the camera plane to $t_{i}^{bg}$ on the background plane, and its distance $d_{i}$ to move from $t_{i}^{c}$ towards $t_{i}^{bg}$ is written as:
$d_{i}=\left|k_{init}\times I_{dep}(\text{center}(M_{i}^{of})\right|.$ (4)
Here, $k_{init}$ represents the depth scaling factor that adjusts the normalized depth values to the appropriate scale. Because the pinhole camera is used to capture the 3D scene, we then raycast from the camera lens $t_{c}$ to $t_{i}^{bg}$ with distance calculated based on $d_{i}$ such that the perpendicular distance moved from the camera plane towards $t_{i}^{bg}$ remains unchanged. The 3D coordinate of object $i$ is:
$t_{i}=t_{c}+\frac{d_{i}}{\left|t_{i}^{bg}-t_{i}^{c}\right|}\times(t_{i}^{bg}-t_{c}).$ (5)
The rotation matrix $R_{i}$ of the 6D pose of object $i$, $P_{i}=[R_{i}\mid t_{i}]$, is set to identity matrix because the reconstructed 3D model $O_{i}$ is aligned in its canonical pose.
Object Scale Calibration   After integrating all 3D object models $\{O_{k}\}_{k=1}^{K}$ into the 3D scene, each with a pose $\{P_{k}=[R_{k}\mid t_{k}]\}_{k=1}^{K}$ and an initial scale $\{S_{k}^{init}\}_{k=1}^{K}$, we refine their scales to accurately reflect depth variations. Through the lens of the pinhole camera with pose $P_{c}$, we render binary masks for each object and evaluate the length of their contour lines relative to their occlusion-free masks. The adjusted, final scale of object $i$ can be expressed as:
$S_{i}^{adj}=S_{i}^{init}\times\frac{\text{arcLength}(\text{findContour}(M_{i}^{of}))}{\text{arcLength}(\text{findContour}(M_{i}^{rend}))},$ (6)
where $M_{i}^{rend}$ is the rendered mask of object $i$.
Finding Object Principle Axes   We determine the principal axes (x-axis, y-axis, and z-axis) of each object, which is essential for our novel 3D motion planning method, explained in Sec. 3.3.1. Using the minimal oriented bounding box (OBB), we encapsulate each object in the smallest possible box, aligning the object’s principal axes with the box’s dimensions. OBB employs principal component analysis on the object’s point cloud to identify directions of maximum, second greatest, and least variance, which naturally correspond to the object’s orientation along its length, width, and height. This approach is straightforward and effective, especially since our input images primarily consist of rigid objects.
3.2.4 Image and Spatial Context Understanding
With accurate camera pose, background plane pose, and object poses and scales, we can reconstruct the 3D scene, denoted as $V_{0}$, from the input image $I$. Before using VLM for motion planning within the 3D scene based on proposed tasks, we conduct in-context learning to familiarize VLM with the reconstructed 3D environment. We feed $V_{0}$ with visualized object axes (see Fig. 2B). Visualizing 3D spatial information is pivotal in improving VLMs’ understanding of 3D spatial contexts derived from 2D images, validated by 3DAxiesPrompts [35]. To augment VLM’s understanding, we also feed a paragraph describing the objects’ poses, sizes, and principal axes in physical units, alongside their spatial relationships. This information is derived entirely from our reconstructed 3D scene. As a concrete example for the left image on Fig. 2:
‘‘Obj 0 spatial context:
- 3D center: [-21.0, 101.0, 4.0] cm
- Local x-axis (towards ‘right’): [0.8637, 0.293, -0.41]
- Local y-axis (towards ‘back_left’): [-0.5943, 0.7879, 0.1614]
- Local z-axis (towards ‘front’): [-0.1606, -0.7822, 0.6019]
Obj 0 size: 29.55 cm x 26.91 cm x 13.19 cm (WxDxH)
Obj 0 - Closest per direction: right: Obj 1 (29.16 cm); down: Obj 4 (15.72 cm) …’’
3.3 Planning and Task Execution within Reconstructed 3D Scene
3.3.1 Axes-Constrained Motion Planning through Waypoints
We introduce a novel method to guide VLM to conduct motion planning within a 3D scene based on a proposed task by planning motion waypoints along the manipulating object’s principal axes. More specifically, we define four types of manipulations that VLM can use:
Rotation type 1: Axial rotation. The object rotates around its principle axes.
Rotation type 2: Rotation relative to the target object.
- pitch: Tilt similar to pouring water, around a horizontal axis formed by the cross product of the connecting directional vector and the target’s vertical axis.
- yaw: Horizontal rotation, like a camera panning, around a vertical axis formed by the cross product of the connecting directional vector and the pitch axis.
- roll: Rotation like drilling a surface, around the connecting directional vector.
Translation type 1: Defines the goal relative to the target object’s principle axes, with translation values for its $[x,y,z]$ axes in centimeters. $[0,0,0]$ cm indicates the goal is the center of the target object.
Translation type 2: Sets the goal relative to a directional vector between two reference objects, specifying how far (in cm) object 1 should move towards or away from object 2 along this vector.
Since VLM inherently lacks the capability to provide 3D coordinates and low-level actions directly [65, 17], our method offers a practical workaround by translating natural language instructions into precise motion waypoints. This approach significantly enhances VLM’s utility in spatial reasoning and manipulation tasks without requiring direct 3D coordinate generation capabilities. Also, the four types of manipulations we defined are both simple and comprehensive, covering a broad spectrum of manipulation tasks.
As a concrete example, given the image on the left of Fig. 2, with ‘‘Task name: Can to Bowl Transfer’’, VLM will plan as follows:
‘‘Task Name: Can to Bowl Transfer
Manipulating obj idx: 3
Interacting obj idx: 4
1. Move Manipulating Obj [3] to [6, 0, 7] cm relative to Target Obj [4]’s local [x, y, z] axes.
2. rotate_wref: Rotate Manipulating Obj [3] relative to Target Obj [4] around [pitch] axis by [75] degrees.’’
3.3.2 Trajectory Generation and Optimization
Trajectory Generation   With the waypoints planned by VLM, we generate the manipulating object’s trajectory using path planning algorithm, specifically rapidly-exploring random tree star (RRT*) [29]. To generate accurate collision-free path, we perform K-means clustering on the point clouds of object 3D model with a high number of clusters, segmenting the object mesh into discrete voxels and treating each voxel as an obstacle. Then, to accurately consider the manipulating object’s dimensions, we grow the size of each voxel by its dimensions.
Handling VLM Planning Discrepancies   The waypoints generated by VLM are typically accurate and practical. Nonetheless, there are instances where the waypoints suggested by VLM lead to collisions as determined by the RRT* planner. This discrepancy is less about VLM’s misunderstanding of the objects’ sizes and their spatial relationship and more about the precision level of the waypoints, which may not match the exacting standards of the RRT* planner’s outcomes. To resolve this, we implement Gaussian sampling around the initially planned waypoints whenever a collision is detected. The sampling strategy is guided by a predefined set of geometric rules. Please see our Appendix. G for details.
Trajectory Optimization   Finally, to ensure our trajectory is natural and smooth, we linearly interpolate rotation and interpolate translation using cubic spline. The final trajectory for proposed task $q\in[1,Q]$ is denoted as $V^{q}=\{V_{r}^{q}\}_{r=0}^{R}$. Each pose in the final trajectory $V^{q}$ is rendered as a frame in the rendering engine, where $V_{0}^{q}=V_{0}$, the rendering of the reconstructed 3D scene.
4 Experiments
Figure 4: Dataset statistics. Our dataset presents 51 scenes—13 from NOCS and 38 captured from varied perspectives—featuring a wide range of object categories, quantities, and a diverse set of tasks, task videos, and planned trajectories.
Implementations   We use ChatGPT-4 [47] as our VLM. The language-guided segmentation model is Language Segment-Anything [43] and the repainting model is LaMa [60]. We use One-2-3-45++ [38] for single-view 3D reconstruction and Depth Anything [69] for depth estimation. For partial 3D scene reconstruction, we use Blender [16] as the 3D software. The pinhole camera focal length and sensor width are set to 80 mm and 36 mm, with $\delta_{init}=2.5e-4,k_{init}=1$. All inference is run on 1 NVIDIA A10 GPU with 24GB RAM.
Dataset   As we are the first to propose zero-shot task hallucination, there lacks a pre-existing dataset for evaluation. Therefore, we craft a diverse evaluation dataset by combining self-captured photos and scenes from the NOCS dataset [62]. We capture 38 photos using an iPhone 12 Pro Max. From NOCS, we use its real-world part of both the training and test sets, which encompasses 13 distinct scenes. Since each scene is a video sequence, we randomly extract one frame as our data. We standardize all image dimensions by resizing all to $640\times 480$.
Dataset Statistics   Our dataset covers diverse scenes (e.g., office, kitchen, bathroom), and features a rich diversity of object categories (116) and quantities (185), with each image containing $1-7$ objects and $1-3$ tasks proposed for each object (278 tasks/task videos/planned trajectories in total). The dataset’s diversity is further enhanced by the variety of perspectives from which the images are captured or selected (e.g., frontal, top-down, side views). This deliberate choice of diverse angles, both in our own image capturing process and through the random extraction of frames from NOCS, aims to simulate a realistic and challenging array of scenes for evaluation. See Fig. 4 for statistics and visuals.
4.1 Task Diversity
Table 1: Comparison of task diversity. We sample 106 proposed tasks for fair comparison with RoboGen and previous RL benchmarks.
Ours RoboGen [66] Behavior-100 [58] RLbench [26] MetaWorld [72] Maniskill2 [22]
Number of Tasks 106 106 100 106 50 20
Self-BLEU $\downarrow$ 0.269 0.284 0.299 0.317 0.322 0.674
Embedding Similarity $\downarrow$ 0.154 0.165 0.210 0.200 0.263 0.194
We evaluate the diversity of the proposed tasks in terms of semantic meaning using Self-BLEU and the embedding similarity [74] following RoboGen [66], where lower scores mean higher diversity. We also compare with previous reinforcement learning (RL) benchmarks. From Table. 1, ours method generates most diverse tasks as our pipeline is open to all scenes with no constraint.
4.2 Qualitative Demonstration
Figure 5: Qualitative results. With diverse input scenes and proposed tasks, our model produces high-quality task execution visualizations with geometric awareness that aligns with the task descriptions. Zoom in for better view.
We present qualitative examples spanning various scenes and tasks, illustrating our pipeline’s capability to produce diverse, visually appealing, and accurate hallucinated task videos for a wide range of input images, as shown in Fig. 5. Additional examples are available at our project page.
Figure 6: Comparison with 3DIT [18] and Runway Gen-2 [52] on rotation and translation tasks. 3DIT fails to maintain consistency and precision across interpolated frames. Runway produces unpredictable artifacts. Our task videos are geometric-aware, photo-realistic with fine-grained and consistent object motion. Zoom in for better view.
We also compare with our baselines, SOTA 3D-aware image editing and video generation models, 3DIT [18] and Runway Gen-2 [52], as shown in Fig. 6. We limit comparisons to rotation and translation tasks to avoid bias against 3DIT and Runway in more nuanced context-dependent tasks (e.g., slicing an apple), where their performances fall short. For 3DIT, we generate and connect frames with our planned trajectories. For Runway, we prompt it with task descriptions.
4.3 Human Preference Evaluation: User Study
Table 2: User study. We ask 25 users to rate (scale $1-5$) and rank (scale $1-3$) for 5 rotation and 5 translation tasks, and rate (scale $1-5$) for 10 complex context-dependent tasks. All results take average. Note that our user study size is similar to those representative works such as ControlNet [73] and Prompt-to-Prompt [23].
Rotation Translation Context-Dependent Rating $\uparrow$ Ranking $\downarrow$ Rating $\uparrow$ Ranking $\downarrow$ Rating $\uparrow$ 3DIT [18] 1.93 2.34 1.90 2.27 - Runway [52] 1.55 2.62 1.43 2.68 - Ours 4.58 1.04 4.43 1.05 4.29
We rely on human preference evaluation as one of our quantitative metrics. We let users rate and rank the rotation and translation task videos generated by 3DIT [18] and Runway Gen-2 [52] against our own in terms of both video quality and task description alignment. For these complex context-dependent tasks, we instead ask users to judge the videos’ authenticity and execution relative to human action, and to encapsulate their perception of the action in our videos with a single sentence. The usage of sentence description will be explained in the subsequent Sec. 4.4. Table. 2 reveals the details and results.
4.4 Understanding Hallucination
Machine Understanding
Table 3: Results for machine understanding (classification) on all 278 task videos.
Machine
Raw Acc $\uparrow$ 0.974
Fal-Pos Rate $\downarrow$ 0.363
True Acc $\uparrow$ 0.611
We assess the interpretability of our generated task videos from a machine’s perspective using SOTA video understanding model, Video-LLaVA-7B [33]. We use two approaches: binary classification and descriptive generation. For classification, we feed the model with the task descriptions generated by VLM and ask question (is the video $\ldots$?). In generation, we prompt Video-LLaVA-7B to articulate its interpretation of the tasks depicted in our videos. To quantify the correspondence between the model’s perception and the tasks, we use OpenCLIP cosine similarity score [14].
However, we find that even SOTA video understanding model shows limited performance. To assess false positive rate in classification, we deliberately misalign the sequence of generated task videos with their corresponding task descriptions, expecting a theoretical accuracy of 0%. Contrary to expectations, Video-LLaVA-7B reports a false positive rate of 36.3%. To adjust for this anomaly, we subtract this rate from the model’s raw accuracy for correctly aligned video-task pairs. This method, while unconventional, provides a more fair and reasonable evaluation of machine video understanding, underscoring the current challenges faced by video understanding models in accurately interpreting complex video content. Results in Table. 3.
Table 4: Results for machine understanding (generation) on all 278 task videos and human understanding, where 25 users write descriptions for 5 videos.
Machine Human
OpenCLIP $\uparrow$ 0.636 0.823
Human Understanding   We aim to understand human perception of our generated task videos by asking participants to provide a one-sentence description. We then evaluate the alignment between these descriptions and the ground truth task descriptions proposed by VLM, using OpenCLIP [14]. Table. 4 reveals a high degree of alignment. Intriguingly, it appears humans understand our hallucinated videos more accurately than machines do. We hypothesize this discrepancy stems from the limitations of current video understanding models, whereas humans draw on their prior experiences for a deeper comprehension.
4.5 Task Proposal Ablation Study
Table 5: Task proposal ablation study. We use all 278 proposed tasks.
Ours Ours w/o obj-init Ours w/o role-play Ours w/o both
Involved Objects Number $\uparrow$ 185 144 185 146
Self-BLEU $\downarrow$ 0.273 0.280 0.276 0.282
Embedding Similarity $\downarrow$ 0.143 0.156 0.141 0.159
We aim to understand if our approach of task proposal, employing role-play and object-based initialization, yields the most diverse and feasible tasks by modifying the prompt to exclude these techniques. In addition to using Self-BLEU and the embedding similarity, we also count the number of objects involved across all 278 proposed tasks.
The results in Table. 5 indicate that our method yields the most diverse tasks. Evaluating on all 278 proposed tasks shows a slight rise in Self-BLEU and a decrease in embedding similarity compared to the 106 tasks sampled (Table. 1). This difference arises because more task descriptions lead to repetitive structures and wording, yet the semantics grow more varied. Excluding object-based initialization reduces diversity by causing the VLM to overlook certain objects. Semantic meaning of tasks cannot reflect feasibility. Instead, we show instances where VLM generates impractical tasks without role-play in Appendix. H. We find that the impact of removing role-play is more complex, affecting task diversity less (Table. 5) but reducing task feasibility.
5 Discussion
We proposed a framework for zero-shot task hallucination. From the observation of a single image, our framework is able to discover diverse tasks and hallucinate task executions as geometric-aware and photo-realistic videos with fine-grained and consistent object motion. We see our work as a step towards equipping machines with the profound human capacity, in which humans can recall past knowledge and anticipate future scenarios through visual perception. Our generated task videos and planned trajectories could spark exciting applications, such as letting robots automatically explore and perform tasks in new environments.
Limitations   Not all generated task videos achieve the highest quality. We identified key factors that contribute to the failure cases. Please see Appendix. I.
Appendix
F Project Page
All code, data, and additional videos are available at our project page: https://dannymcy.github.io/zeroshot_task_hallucination/.
G Additional Details of Method
G.1 Matching-Based Camera 6D Pose Estimation
For single-view 3D reconstruction models which do not provide the 6D pose, $P_{c}=[R_{c}\mid t_{c}^{origin}]$, of the pinhole camera, we develop an efficient method to determine the pose by comparing object masks with rendered 3D model templates, inspired from matching-based 6D pose estimation works [34, 46, 9, 41].
We generate a set of object templates, denoted as $\{T_{j}^{i}\}_{j=1}^{J}$, each rendered from the object’s 3D model $O_{i}$. These templates are created by positioning the camera at various locations on an icosphere surrounding the object in $SE(3)$ space, which simulates a spherical coverage around the object to capture its geometry from all angles uniformly. For each template $T_{j}^{i}$, we compute a matching score against the occlusion-free object mask $M_{i}^{of}$.
We propose a simple yet effective score matching method. We draw a bounding rectangle around the segmented object inside $M_{i}^{of}$ and across all $\{T_{j}^{i}\}_{j=1}^{J}$, and crop the bounding rectangle. We then calculate the shape similarity between the contour line of cropped $M_{i}^{of}$ and that of each cropped $T_{j}^{i}$ using Hu moments [25]. Additionally, we crop and resize the bounding rectangle to the same dimension, and evaluate the similarity based on the pixel area of the cropped and resized masks. Our score matching method can be formulated as:
$m_{i}^{A,h}&=\text{Hu}(\text{findContour}(\text{crop}(M_{i}^{of}))),\;\;\;\;\;\;pa_{i}^{A}=\text{sum}(\text{resize}(\text{crop}(M_{i}^{of})))\\ m_{i,j}^{B,h}&=\text{Hu}(\text{findContour}(\text{crop}(T_{j}^{i}))),\;\;\;\;\;\;\;\;pa_{i,j}^{B}=\text{sum}(\text{resize}(\text{crop}(T_{j}^{i})))\\ \mathcal{L}(M_{i}^{of},T_{j}^{i})&=\alpha\left|1-\frac{\min(pa_{i}^{A},pa_{i,j}^{B})}{\max(pa_{i}^{A},pa_{i,j}^{B})}\right|+\beta\sum_{h=1}^{7}\left|\frac{1}{\text{sgn}(m_{i}^{A,h})\cdot\log(m_{i}^{A,h})}-\frac{1}{\text{sgn}(m_{i,j}^{B,h})\cdot\log(m_{i,j}^{B,h})}\right|\end{aligned}$
This dual approach allows for a comprehensive comparison that incorporates both the geometric configuration and the scale of the object representations. The best-matched template can be found by $\arg\min_{j=1}^{J}\mathcal{L}(M_{i}^{of},T_{j}^{i})$, and we set its corresponding camera pose to be the 6D pose of our pinhole camera, denoted as $P_{c}=[R_{c}\mid t_{c}^{origin}]$.
The camera now captures the object’s canonical view, centered at the 3D origin. Subsequently, we adjust the camera’s translation component to focus on the center of the inpainted background $I_{bg}$, resulting in $P_{c}=[R_{c}\mid t_{c}]$, as detailed in our main paper.
G.2 Sampling Rules for VLM Planning Discrepancies
Figure 7: Feasibility comparison of the proposed tasks with and without role-play.
We define a set of geometric rules for Gaussian sampling around the initially planned waypoints whenever a collision is detected. In our 3D coordinate system, positive x-axis [1, 0, 0] points right, positive y-axis [0, 1, 0] is away from viewer, positive z-axis [0, 0, 1] is up. For translation type 1, we denote the goal pose relative to the target object’s principle axes as [dx, dy, dz]. For translation type 2, we denote the distance that object 1 moves towards object 2 as dD. The set of geometric rules are as follows:
if type 1 &\&& [dx, dy, dz] = [0, 0, 0]: sample along [x, y, z] axes
elif type 1 &\&& dx = 0 &\&& dy = 0 &\&& dz != 0: sample along [z] axis, $\text{z}_{\text{sampled}}\cdot$ dz > 1
elif type 1 &\&& dz = 0: sample along [x, y] axes
elif type 1 &\&& dz != 0: sample along [x, y, z] axes, $\text{z}_{\text{sampled}}\cdot$ dz > 1
if type 2: sample along the connecting directional vector, $\text{D}_{\text{sampled}}\cdot$ dD > 1
G.3 Hand Insertion
Each pose in the final trajectory $V^{q}$ is rendered as a frame in the rendering engine. To enhance human comprehension and perception, we integrate a 3D model of “a magic hand” [68] in proximity to the object being manipulated.
H Task Proposal Comparison of Feasibility
We show instances where ChatGPT-4 [47] generates impractical tasks without role-play in Fig. 7. These unfeasible tasks usually cannot be performed by a robotic/human hand due to physical limitations or require additional tools beyond the capabilities of the hand alone.
I Failure Cases
Figure 8: Key factors causing failure cases. For each factor, the leftmost image is the input.
Not all generated task videos achieve the highest quality. We identify several key factors that contribute to the failure cases. Fig. 8 shows visualization.
Segmentation Failure   When an image contains objects with very similar appearances, ChatGPT-4 can distinguish them, providing unique short descriptions. For instance, in the input image on the left of Fig. 8(a), ChatGPT-4 differentiates two white bowls as ‘‘Object 2: bowl of color white with blue pattern, ‘‘Object 3: bowl of color white with black pattern". Nonetheless, the language-guided segmentation model, Language Segment-Anything [43], may fail to capture the nuance, incorrectly merging object masks. This issue is depicted in Fig. 8(a)’s right reconstructed scene, where Language Segment-Anything recognizes and merges the two distinct bowls as one.
Inpainting Failure   In rare cases, the repainting model LaMa [60] introduces unintended elements into and distorts object masks, illustrated by the deformed white bowl in Fig. 8(b)’s right reconstructed scene. Such distortions result in inaccurately reconstructed 3D models.
3D Model Reconstruction Failure   The model One-2-3-45++[38], despite its strength, occasionally reconstructs less-than-ideal 3D object models, as illustrated by the misshapen lamp shade in Fig. 8(c)’s right reconstructed scene.
Motion Planning Failure   ChatGPT-4 occasionally produces suboptimal planned motions, primarily due to incorrect use of object principal axes and inaccurate translation measurements. For instance, as shown in Fig. 8(d), one would expect the spoon to be inserted into the mug. However, ChatGPT-4’s planning positions the spoon too high, making it appear to float above the mug.
J Prompt Details
Figure 9: Example input scene. (a) the original input image, (b) the reconstructed scene with visualized axes.
We show exact prompts and example responses from ChatGPT-4 for the white plate with coffee grounds shown in the input images of Fig. 9.
Interactive Objects Identification Prompt.
Input:
RGB image (640, 480) = (width, height) with multiple objects.
Your task is to identify and name useful objects (in term of affordances) with precise color and texture. Exclude background objects. Do not use vague phrase like multi-colored.
Please write in the following format. Do not output anything else:
Object idx (actual integer, start from 0): x of color y
An example response from ChatGPT-4:
‘‘Object 0: sink of color silver
Object 1: faucet of color silver
Object 2: dish soap of color yellow
Object 3: hand soap of color purple
Object 4: sponge of color blue
Object 5: white plate with coffee grounds of color brown’’
Task Proposal Prompt.
Input:
1. RGB image (640, 480) = (width, height) with multiple objects.
2. Detected objects with index.
You are a single robot hand working in this image scene to perform simple household tasks. Tasks must be discovered from the image. Consider objects’ affordances and feel free to make assumptions (e.g., a bowl can contain water) and interactions with other objects (e.g., pouring water from a cup into a bowl).
Task types:
1. Interaction between the manipulating object and one of the detected objects (involve translation, or translation + rotation).
2. Rotate manipulating object (involve rotation).
Strictly follow constraints:
1. Exclude tasks involving assembly or disassembly of objects.
2. Exclude tasks involving cleaning or functionality testing.
3. Exclude tasks involving imaginary objects.
4. Manipulating object moves; interacting object static.
5. Assume all objects are rigid, without joints or moveable parts (i.e., cannot deform, disassemble, transform). This applies even to objects that are typically articulated (e.g., laptop).
Propose 3 tasks (2 interaction, 1 rotation) for manipulating Object 5. Write in the following format. Do not output anything else:
Task Name: xxx
Manipulating obj idx: 5
Interacting obj idx: obj_idx (actual integer, or manipulating obj idx)
Description: basic descriptions.
An example response from ChatGPT-4:
‘‘Task Name: Plate Orientation
Manipulating obj idx: 5
Interacting obj idx: 5
Description: Rotate the plate 180 degrees horizontally to change its orientation while remaining in the same location on the sink.
Task Name: Plate Relocation
Manipulating obj idx: 5
Interacting obj idx: 0
Description: Move the plate from its current location to the right side of the sink to clear space for other kitchen activities.
Task Name: Washing Plate
Manipulating obj idx: 5
Interacting obj idx: 1
Description: Position the plate beneath the faucet and simulate the process of rinsing it.’’
Image and Spatial Context Understanding Prompt.
Inputs:
1. RGB image (640, 480) = (width, height) with multiple objects and their visualized local axes (x red, y green, z blue).
2. Detected objects with index.
3. For each detected object, its 3D center, local xyz-axes, size, and spatial relationship relative to other objects.
The 3D coordinate system of the image is in centimeters and follows Blender. Positive x-axis [1, 0, 0] right, positive y-axis [0, 1, 0] away from viewer, positive z-axis [0, 0, 1] up. Positive rotation is counter-clockwise around all axes.
Your task is to learn the spatial context. Also, consider how a robot or human hand might navigate or interact with these objects, focusing on proximity, accessibility, and collision avoidance. Do not output.
An example input of object spatial information to ChatGPT-4:
‘‘Obj 0 spatial context:
3D center: [-21.0, 90.0, -14.0] cm
Local x-axis (towards ’right’): [0.9932, -0.0882, -0.0766]
Local y-axis (towards ’back’): [0.1104, 0.9234, 0.3675]
Local z-axis (towards ’up’): [0.0384, -0.3735, 0.9268]
Obj 1 spatial context:
3D center: [-35.0, 85.0, 8.0] cm
Local x-axis (towards ’front_right’): [0.8565, -0.5126, 0.0612]
Local y-axis (towards ’back_right’): [0.5151, 0.8563, -0.0377]
Local z-axis (towards ’up’): [-0.0331, 0.0638, 0.9974]
Obj 2 spatial context:
3D center: [31.0, 80.0, 13.0] cm
Local x-axis (towards ’right’): [0.9971, -0.0062, 0.0756]
Local y-axis (towards ’back’): [0.008, 0.9997, -0.0226]
Local z-axis (towards ’up’): [-0.0754, 0.0231, 0.9969]
Obj 3 spatial context:
3D center: [33.0, 71.0, 8.0] cm
Local x-axis (towards ’right’): [0.8162, -0.0206, 0.5775]
Local y-axis (towards ’back’): [-0.0169, 0.9981, 0.0595]
Local z-axis (towards ’up’): [-0.5776, -0.0583, 0.8142]
Obj 4 spatial context:
3D center: [22.0, 97.0, 3.0] cm
Local x-axis (towards ’right’): [0.999, 0.0383, -0.0231]
Local y-axis (towards ’back’): [-0.0399, 0.9966, -0.0721]
Local z-axis (towards ’up’): [0.0203, 0.073, 0.9971]
Obj 5 spatial context:
3D center: [-12.0, 78.0, -13.0] cm
Local x-axis (towards ’right’): [0.9998, -0.0216, -0.001]
Local y-axis (towards ’back’): [0.0216, 0.9998, 0.0052]
Local z-axis (towards ’up’): [0.0009, -0.0052, 1.0]
Obj 0 size: 47.63 cm x 39.15 cm x 22.25 cm (WxDxH)
Obj 1 size: 19.11 cm x 13.82 cm x 17.83 cm (WxDxH)
Obj 2 size: 9.69 cm x 9.19 cm x 15.63 cm (WxDxH)
Obj 3 size: 7.74 cm x 6.91 cm x 13.97 cm (WxDxH)
Obj 4 size: 9.57 cm x 2.17 cm x 5.35 cm (WxDxH)
Obj 5 size: 12.17 cm x 11.90 cm x 7.03 cm (WxDxH)
Obj 0 - Closest per direction: up: Obj 1 (26.75 cm); right: Obj 4 (46.28 cm); front_right: Obj 5 (14.55 cm)
Obj 1 - Closest per direction: down: Obj 0 (26.75 cm); right: Obj 5 (31.93 cm)
Obj 2 - Closest per direction: left: Obj 5 (50.92 cm); front: Obj 3 (11.06 cm); back_left: Obj 4 (21.93 cm)
Obj 3 - Closest per direction: left: Obj 5 (49.76 cm); back: Obj 2 (11.06 cm)
Obj 4 - Closest per direction: left: Obj 0 (46.28 cm); front_right: Obj 2 (21.93 cm); front: Obj 3 (28.77 cm); front_left: Obj 5 (41.97 cm)
Obj 5 - Closest per direction: back_left: Obj 0 (14.55 cm); left: Obj 1 (31.93 cm); right: Obj 3 (49.76 cm); back_right: Obj 4 (41.97 cm)’’
Motion Planning Prompt.
Inputs:
1. RGB image (640, 480) = (width, height) with multiple objects and their visualized local axes (x red, y green, z blue).
2. Detected objects with index.
3. Simple household tasks and descriptions to be performed by a single robot hand.
Your goal is to plan fine-grained motions for the manipulating object to complete the tasks using four manipulations, explained as follows:
Rotation:
rotate_self: Axial rotation. The object rotates around its local [x/y/z] axis by [degrees].
rotate_wref: Rotation relative to the target object:
- pitch: Tilt similar to pouring water, around a horizontal axis formed by the cross product of the connecting directional vector and the target’s z-axis.
- yaw: Horizontal rotation, like a camera panning, around a vertical axis formed by the cross product of the connecting directional vector and the pitch axis.
- roll: Rotation like a drill entering a surface, around the connecting directional vector.
The degrees can be specified in two ways:
- Exact [degrees]. Positive values rotate the manipulating object towards the target object.
- Fixed_towards/fixed_back. ’fixed_towards’ orients the object towards the target, mimicking actions like pouring (pitch), facing (yaw), or drilling into (yaw+roll) the target. ’fixed_back’ reverses this alignment.
Translation:
translate_tar_obj: Defines the goal relative to the target object’s local axes, with translation values for its [local_x, local_y, local_z] axes in centimeters. [0, 0, 0] cm indicates the goal is the center of the target object.
translate_direc_axis: Sets the goal relative to a directional vector between two reference objects, specifying how far (in cm) object 1 should move towards or away from object 2 along this vector (positive closer, negative away). Object indices must differ, and if one reference object is the manipulating object, its current location is used.
Strictly follow caveats:
1. Apply rotate_wref thoughtfully and sequentially around different axes as needed.
2. Use the provided spatial information and image effectively for understanding and planning within the 3D scene.
3. Combine common physical understanding with the scene’s spatial details (like relative positions and sizes of objects) for strategic planning.
4. Remember that objects’ local axes’ positive directions might require using negative values in rotation and translation for authentic motion planning.
Plan as below. Fill in obj_idx based on the tasks.
rotate_self: Rotate Manipulating Object [obj_idx] around its local axis [x/y/z] by [degrees].
rotate_wref: Rotate Manipulating Object [obj_idx] relative to Target Object [target_obj_idx] around [pitch/yaw/roll] axis by [degrees/fixed_towards/fixed_back].
translate_tar_obj: Move Manipulating Object [obj_idx] to [a, b, c] cm relative to Target Object [target_obj_idx]’s local [x, y, z] axes.
translate_direc_axis: Move Manipulating Object [obj_idx] [a] cm along the directional vector from Reference Object [ref_obj_1_idx] to Reference Object [ref_obj_2_idx].
Here are some full examples. Please write in the following format. Do not output anything else:
Task Category: Bear rotation
Description: Rotate the toy bear 90 degrees on its vertical axis.
Motion Planning:
Manipulating obj idx: bear_idx (actual integer)
Interacting obj idx: bear_idx (actual integer)
1. rotate_self: Rotate Manipulating Object [bear_idx] around its local axis [z] by [90] degrees.
Task Name: Cup content transfer
Description: Pick up the mug and pour its contents into the bowl.
Motion Planning:
Manipulating obj idx: cup_idx (actual integer)
Interacting obj idx: bowl_idx (actual integer)
1. translate_tar_obj: Move Manipulating Object [cup_idx] to [5, -7, 5] cm relative to Target Object [bowl_idx]’s local [x, y, z] axes.
2. rotate_wref: Rotate Manipulating Object [obj_idx] relative to Target Object [bowl_obj_idx] around [pitch] axis by [fixed_towards].
Task Name: Screwdriver penetration
Description: Use a screwdriver to penetrate an avocado.
Motion Planning:
Manipulating obj idx: screw_idx (actual integer)
Interacting obj idx: avocado_idx (actual integer)
1. translate_tar_obj: Move Manipulating Object [screw_idx] to [-5, -5, 0] cm relative to Target Object [avocado_idx]’s local [x, y, z] axes.
2. rotate_wref: Rotate Manipulating Object [screw_idx] relative to Target Object [avocado_idx] around [yaw] axis by [fixed_towards].
3. rotate_wref: Rotate Manipulating Object [screw_idx] relative to Target Object [avocado_idx] around [roll] axis by [360] degrees.
An example response from ChatGPT-4:
‘‘Task Name: Plate Orientation
Description: Rotate the plate 180 degrees horizontally to change its orientation while remaining in the same location on the countertop.
Motion Planning:
Manipulating obj idx: 5
Interacting obj idx: 5
1. rotate_self: Rotate Manipulating Object 5 around its local axis [z] by [180] degrees.
Task Name: Plate Relocation
Description: Move the plate from its current location to the right side of the sink to clear space for other kitchen activities.
Motion Planning:
Manipulating obj idx: 5
Interacting obj idx: 0
1. translate_tar_obj: Move Manipulating Object 5 to [26.5, 0, 0] cm relative to Target Object 0’s local [x, y, z] axes.
Task Name: Washing Plate
Description: Position the plate beneath the faucet and simulate the process of rinsing it.
Motion Planning:
Manipulating obj idx: 5
Interacting obj idx: 1
1. translate_tar_obj: Move Manipulating Object 5 to [6.5, 0, -5] cm relative to Target Object 1’s local [x, y, z] axes.
2. rotate_self: Rotate Manipulating Object 5 around its local axis [x] by [-15] degrees.
3. rotate_self: Rotate Manipulating Object 5 around its local axis [y] by [-15] degrees.
4. rotate_self: Rotate Manipulating Object 5 around its local axis [y] by [10] degrees.
5. rotate_self: Rotate Manipulating Object 5 around its local axis [x] by [10] degrees.’’